OpenStudy (anonymous):

Integrate ( x^2+1/x ) dx from 1 to e

OpenStudy (anonymous):

\[\int\limits_{1}^{e} (x^2+1) / x dx\]

OpenStudy (anonymous):

Write firs \[ \frac {x^2 +1}x = x + \frac 1 x \] Can you finish it from here?

OpenStudy (anonymous):

first

OpenStudy (anonymous):

so do i integrate them separetly. so integral of x will be x^2/2 and integral of 1/x is ln(x)

OpenStudy (anonymous):

wait but how did u get x + 1/x first